PHOTOFRAGMENTATION DYNAMICS OF ICN^{−}(CO_{2})_{n}

\underline{JOSHUA P. MARTIN}, JOSHUA P. DARR, W. CARL LINEBERGER, JILA and Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309; ANNE B. MCCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

We report photofragmentation studies of mass-selected ICN^{−}(CO_{2})_{n} (n=0-18) cluster anions following 400, 500 and 600 nm excitation. Photodissociation of ICN^{−} following excitation via a 500 nm photon results exclusively in an I^{−} anionic photoproduct. However, excitation at either 400 nm or 600 nm results in 30% CN^{−} and 70% I^{−} anionic photoproducts. Complexation of ICN^{−} with two or more CO_{2} molecules opens a third product channel, recombination of the chromophore on the ground state, a channel that is observable at all three photon energies. The product branching ratios for photodissociation of ICN^{−}(CO_{2})_{n} show sharp contrast to trends previously observed in dihalide studies of IBr^{−}(CO_{2})_{n} and I_{2}^{−}(CO_{2})_{n}, but are strikingly similar to results found for ICl^{−}(CO_{2})_{n}. Notably, a peak in the percentage of recombined photoproducts observed as a function of the number of solvent molecules shifts to larger n values as photon energy increases and does not reach 100% in the first solvation shell. It is likely that an intracluster reaction involving formation of the molecule NCCO_{2}^{−} plays a significant role in the dynamics observed following excitation of the ICN^{−}(CO_{2})_{n} anion. Supported by NSF and AFOSR